# Beyond Flatland: A Geometric Take on Matching Methods for Treatment Effect Estimation

### v1

@@ 1 Introduction
We propose GeoMatching, a framework to estimate treatment effects accounting for the underlying geometry and uncertainty of the data. Instead of working in the space of raw input characteristics, we learn a low-dim latent representation together with a latent Riemannian metric (a generalization of the Euclidean metric) such that raw distances along the original manifold embedded in the high-dim input space are preserved in the low-dim latent space (see Figure 1C). We can then leverage those geometry-aware distances for matching to estimate treatment effects more accurately, as shown in Figure 1D. The latent representation is a Riemannian manifold, where distance between two data points is given by minimum path length along the data manifold. To our knowledge, this work is the first one looking at geometry-aware methods (in a differential geometric sense) for causal inference.
@@ 3.1 Assumptions
We assume that the treatment effect varies smoothly along the manifold of confounders. Back to the motivational example in Section 1, we expect athletes with similar anatomical constraints (reflected by the manifold structure) to respond similarly to different training strategies. As another example, a chemist might want to assess the effect of adding a catalyst to a chemical reaction (treatment) on the chemical reaction rate (outcome) given experimental conditions such as pH, purity of reactants, pressure, temperature, etc (confounders). As we move along the space of possible experimental conditions (manifold), we expect a smooth variation of the treatment effect.
@@ 3.2 A Geometric Take on Treatment Effect Matching
See Appendix 7.1 for further details and precise mathematical definitions of key Riemannian concepts.
@@ 3.3 Algorithm Description
where $\mathrm{vec}\,\mathbf{G}$ stacks the columns of $\mathbf{G}$ and $\otimes$ denotes the Kronecker product. Based on Picard-Lindelöf theorem (Tenenbaum and Pollard, 1985), geodesics are guaranteed to exist and are locally unique given a starting point and initial velocity $\gamma^{\prime}$. See Appendix 7.5 for further details and discussion on computational cost of GeoMatching.
@@ 4 Related Work
Several previous works learn low-dim representations before matching to mitigate undesirable effects of high dimensions in causal inference. Clivio et al. (2022) leverage the representation of intermediate layers of a neural network that predicts the propensity score $p(T|X)$, showing that such representation is approximately a balancing score, i.e., $X\perp T|Z$. Luo and Zhu (2017) and Zhao et al. (2022) propose matching methods on linear projections of the data based on the sufficient dimensionality reduction property, i.e., $X\perp Y|Z$. Wang et al. (2021) and Li et al. (2016) propose matching algorithms for high-dim data: the former relies on variable selection for categorical data; the later conducts and aggregates matching on random projections of the data. Clivio et al. (2023) characterize the bias added to a weighting estimator when using a data representation and propose an objective to learn suitable representations for causal inference. None of these methods account for the geometry of confounders, which is the main goal of this work.
@@
While there has been extensive work on learning suitable latent representations for causal inference, very few have considered the geometry of the latent space for treatment effect estimation. Yan et al. (2024) is the only work we found that proposes to consider the geometry of confounders to improve TE estimation. They propose an optimal transport framework to promote the assignment of low weights for outliers, but they still assume Euclidean distances as opposed to a generic Riemannian distance along the data manifold. To our knowledge, our work is the first one to propose geometry-aware treatment effect estimation based on Riemannian geometry.
@@ 5.1 Experimental Setup
For evaluation, we qualitatively inspect the matched units (geodesic curves) in the latent space, and quantitatively assess performance of treatment effect estimation via Average Treatment Effect (ATE) absolute error, defined as the absolute difference between the true and estimated ATEs. We also report Precision in Estimation of Heterogeneous Effect (PEHE) and distributions of Individual Treatment Effect (ITE) in Appendix 7.3. When plotting the latent representation of GeoMatching, we visualize the magnification factor – also called Riemannian volume – which measures the magnitude of the local distortion of space at location $p\in\mathcal{M}$, i.e., $V_{\mathbf{G}}(p):=\sqrt{\mathrm{det}\,\mathbf{G}(p)}$. Curves that traverse regions with high magnification factor tend to have larger lengths; it can also be seen as how much energy is needed to traverse such space. Geodesics or shortest paths tend to avoid regions with high values of magnification factor, which reflect high uncertainty away from the observed data manifold.
@@ 5.2 Synthetic Swissroll Dataset
We adapt the classic swissroll dataset (see Appendix 7.2 for precise details on the data generation process) for the task of treatment effect estimation. We generate $N=200$ observations where (pre-treatment) covariates $X$ follow a 3D-swissroll manifold, treatment assignment variables $T$ are Bernoulli draws from a smoothly varying sigmoid function along the swissroll manifold, and outcome variables $Y$ are drawn from linear models. To assess the impact of input dimensionality on the different matching strategies, we embed the three-dimensional swissroll in a space of higher dimensionality $D\in\{3,5,10,15,25,50,100\}$, by adding additional noisy (irrelevant) input dimensions.
@@
When looking at performance for treatment effect estimation (Figures 2(d) and 2(e)), GeoMatching yields the most accurate ATEs. As we increase the dimensionality $D$ of input space $\mathcal{X}$, performance of euclidean matching in $\mathcal{X}$ degrades until reaching the same performance as random matching. Unsurprisingly, euclidean matching in $\mathcal{Z}$ remains competitive as $D$ increases since it removes irrelevant dimensions (like GeoMatching) when projecting the data to a low-dim latent space.
@@
(d) Treatment Effect Estimation Performance
@@
Figure 2: Results for swissroll synthetic toy dataset. (a-c) Matched pairs according to baselines and our method GeoMatching, (d) ATE Absolute Errors averaged over 20 random seeds.
@@ 5.3 Semi-synthetic Human Motion Capture
We use a synthetic scenario of human motion capture data, which consists of real-world input covariates and synthetic treatment and outcome assignment variables. The goal of this experiment is to demonstrate i) effectiveness of GeoMatching handling high-dimensional realistic covariates, and ii) robustness of GeoMatching against outlier confounders, in contrast to geometry-agnostic methods.
@@
We adapt the CMU mocap data for the downstream task of treatment effect estimation. To motivate this new setup, consider Double Dutch, a jump rope game that started in the streets and has now advanced to competitions, since 1974.444https://youtu.be/iiEzf3J4iFk?feature=shared&t=49. To play, two people turn two long ropes in opposite directions while, multiple players jump simultaneously, entering the rope area one at a time. Let $X$ be the motion capture of the current player jumping, $T$ the strategy of entrance, either slow ($T=0$) or fast ($T=1$), and $Y$ the probability of a successful entrance, when ropes do not trip up the new jumper’s feet. To make the data more challenging, we intentionally perturb some time frames to simulate outlier confounders in the dataset (i.e., players jumping in a weird, unnatural manner, or motion capture sensors being defective). See Section 7.2 for further details about the data generation process.
@@
Figure 3: Results for semi-synthetic motion capture data. (a) Visualization of the latent space and corresponding high-dimensional confounders. (b-d) Matched pairs according to baselines and our method GeoMatching, (d) Average Treatment Effect Absolute Errors (avg across 20 random seeds).
Figure 3(a) shows a visualization of the data in a 2D-latent space using PCA dimensionality reduction, which preserves the geometric/periodic structure of the data. Figures 3(b) and 3(c) show matched units (geodesic curves) when pairing cases and controls according to Euclidean (baseline) or Riemannian metric (GeoMatching) in the latent space. All baselines are sensitive to the introduced outliers because they do not take into account data uncertainty; in contrast, GeoMatching is robust against the impact of outliers, which results in better ATE estimation, as shown in Figure 3(d). These results align with our a priori expectations, as outliers tend to lie in low-density (high-volume) regions which are unlikely to be traversed by geodesics.
@@ 5.4 Real-world Causality Benchmark Datasets
We evaluate GeoMatching on two real-world datasets. We demonstrate that i) GeoMatching is competitive in more realistic scenarios, and ii) GeoMatching can benefit from semi-supervised settings, where we have access to several unlabelled covariate data, from which only a small subset has treatment and outcome information available. We consider the Infant Health Development Program Dataset and the Lalonde [^5] https://users.nber.org/~rdehejia/nswdata2.html dataset, see Appendix 7.2 for further details.
Table 2(b) summarizes the results on real-world datasets. For PCA, GeoMatching improves upon all baselines. We also include results for the Isomap projection. In this case, Euclidean matching in $\mathcal{Z}$ is the best approach, slightly above GeoMatching. We hypothesize that this is due to a coarse grid when optimizing the $\sigma$ parameter of the Riemannian metric (only 10 categorical values were explored in a grid search setup, see Appendix for details). In the case of the LaLonde dataset, using a Riemannian metric yield improvements in the Original covariate space, as well as in the PCA and NSM latent representations.

### v2

@@ Abstract
keywords:
Causal inference, matching, Riemannian geometry, manifold.
@@ 1 Introduction
We propose GeoMatching, a framework to estimate treatment effects accounting for the underlying geometry and uncertainty of the data. Instead of working in the space of raw input characteristics, we learn a low-dim latent representation together with a latent Riemannian metric (a generalization of the Euclidean metric) such that raw distances along the original manifold embedded in the high-dim input space are preserved in the low-dim latent space (see Figure 1C). We can then leverage those geometry-aware distances for matching to estimate treatment effects more accurately, as shown in Figure 1D. [^1] Figure 1D shows ITE/ATE estimation results corresponding to the synthetic 3D-swissroll dataset, which are later expanded in Section 5.2. Appendix B describes the data generation process in detail. The latent representation is a Riemannian manifold, where distance between two data points is given by minimum path length along the data manifold. To our knowledge, this work is the first one looking at geometry-aware methods (in a differential geometric sense) for causal inference.
@@ 3.1 Assumptions
We assume that the treatment effect varies continuously along the manifold of confounders. Continuity ensures that points close to each other in the latent manifold will have similar treatment effects. This property is central in GeoMatching, as it allows us to ensure that small changes in the input (manifold points) do not lead to large, abrupt changes in the treatment effect. Back to the motivational example in Section 1, we expect athletes with similar anatomical constraints (reflected by the manifold structure) to respond similarly to different training strategies. As another example, a chemist might want to assess the effect of adding a catalyst to a chemical reaction (treatment) on the chemical reaction rate (outcome) given experimental conditions such as pH, purity of reactants, pressure, temperature, etc (confounders). As we move along the space of possible experimental conditions, we expect a smooth variation of the treatment effect.
@@ 3.2 A Geometric Take on Treatment Effect Matching
See Appendix A for further details and precise mathematical definitions of key Riemannian concepts.
@@ 3.3 Algorithm Description
where $\mathrm{vec}\,\mathbf{G}$ stacks the columns of $\mathbf{G}$ and $\otimes$ denotes the Kronecker product. Based on Picard-Lindelöf theorem (Tenenbaum and Pollard, 1985), geodesics are guaranteed to exist and are locally unique given a starting point and initial velocity $\gamma^{\prime}$. See Appendix E for further details and discussion on computational cost of GeoMatching.
@@ 4 Related Work
Several previous works learn low-dim representations before matching to mitigate undesirable effects of high dimensions in causal inference. Clivio et al. (2022) leverage the representation of intermediate layers of a neural network that predicts the propensity score $p(T|X)$, showing that such representation is approximately a balancing score, i.e., $X\perp T|Z$. Luo and Zhu (2017) and Zhao et al. (2022) propose matching methods on linear projections of the data based on the sufficient dimensionality reduction property, i.e., $X\perp Y|Z$. Wang et al. (2021) and Li et al. (2016) propose matching algorithms for high-dim data: the former relies on variable selection for categorical data; the later conducts and aggregates matching on random projections of the data. Clivio et al. (2023) characterize the bias added to a weighting estimator when using a data representation and propose an objective to learn suitable representations for causal inference. In contrast to our work, none of these methods account for the geometry of confounders.
@@
While there has been extensive work on learning suitable latent representations for causal inference, very few have considered the geometry of the latent space for treatment effect estimation. Yan et al. (2024) is the only work we found that proposes to consider the geometry of confounders to improve TE estimation. They propose an optimal transport framework to promote the assignment of low weights for outliers, but they still assume Euclidean distances as opposed to a generic Riemannian distance along the data manifold. To our knowledge, our work is the first one to propose geometry-aware TE estimation based on Riemannian geometry.
@@ 5.1 Experimental Setup
For evaluation, we qualitatively inspect the matched units (geodesic curves) in the latent space, and quantitatively assess performance of treatment effect estimation via Average Treatment Effect (ATE) absolute error, defined as the absolute difference between the true and estimated ATEs. We also report Precision in Estimation of Heterogeneous Effect (PEHE) and distributions of Individual Treatment Effect (ITE) in Appendix C. When plotting the latent representation of GeoMatching, we visualize the magnification factor – also called Riemannian volume – which measures the magnitude of the local distortion of space at location $p\in\mathcal{M}$, i.e., $V_{\mathbf{G}}(p):=\sqrt{\mathrm{det}\,\mathbf{G}(p)}$. Curves that traverse regions with high magnification factor tend to have larger lengths; it can also be seen as thr energy needed to traverse such space. Geodesics or shortest paths tend to avoid regions with high values of magnification factor, which reflect high uncertainty away from the observed data manifold.
@@ 5.2 Synthetic Swissroll Dataset
We adapt the classic swissroll dataset (see Appendix B for precise details on the data generation process) for the task of treatment effect estimation. We generate $N=200$ observations where (pre-treatment) covariates $X$ follow a 3D-swissroll manifold, treatment assignment variables $T$ are Bernoulli draws from a smoothly varying sigmoid function along the swissroll manifold, and outcome variables $Y$ are drawn from linear models. To assess the impact of input dimensionality on the different matching strategies, we embed the 3D-swissroll in a space of higher dimensionality $D\in\{3,5,10,15,25,50,100\}$, by adding additional noisy (irrelevant) input dimensions.
@@
Looking at performance for TE estimation in Figures 2(d) and 2(e), GeoMatching yields the most accurate ATEs. As we increase the dimensionality $D$ of input space $\mathcal{X}$, performance of euclidean matching in $\mathcal{X}$ degrades until reaching the same performance as random matching. Unsurprisingly, euclidean matching in $\mathcal{Z}$ remains competitive as $D$ increases since it removes irrelevant dimensions (like GeoMatching) when projecting the data to a low-dim latent space.
@@
(d) TE Estimation Performance
@@
Figure 2: Results for swissroll synthetic toy dataset. (a-c) Matched pairs according to baselines and our method GeoMatching, (d-e) ATE Absolute Errors averaged over 20 random seeds.
@@ 5.3 Semi-synthetic Human Motion Capture
We use a semi-synthetic scenario of human motion capture data, which consists of real-world input covariates and synthetic treatment and outcome assignment variables. Such setup lies in-between an ideal/simplified synthetic scenario (like the 3D-swissroll example from previous section) and a fully realistic scenario where there is no access to ground truth ITEs (like the LaLonde dataset in Section 5.4). The goal of this experiment is to demonstrate i) effectiveness of GeoMatching handling high-dim realistic covariates, and ii) robustness of GeoMatching against outlier confounders, in contrast to geometry-agnostic methods.
@@
We adapt the CMU mocap data for the downstream task of TE estimation. To motivate this new setup, consider Double Dutch, a jump rope game that started in the streets and has now advanced to competitions, since 1974.555https://youtu.be/iiEzf3J4iFk?feature=shared&t=49. To play, two people turn two long ropes in opposite directions while, multiple players jump simultaneously, entering the rope area one at a time. Let $X$ be the motion capture of the current player jumping, $T$ the strategy of entrance, either slow ($T=0$) or fast ($T=1$), and $Y$ the probability of a successful entrance, when ropes do not trip up the new jumper’s feet. To make the data more challenging, we intentionally perturb some time frames to simulate outlier confounders in the dataset (i.e., players jumping in a weird, unnatural manner, or motion capture sensors being defective). See Appendix B for further details about the data generation process.
@@
Figure 3: Results for semi-synthetic motion capture data. (a) Visualization of the latent space and corresponding high-dim confounders after PCA projection. Outliers lie outside the intrinsic data manifold. (b-c) Matched units (geodesic curves) when pairing cases and controls according to Euclidean (baseline) or Riemannian metric (GeoMatching) in the latent space. (d) Average Treatment Effect Absolute Errors (avg across 20 random seeds).
Figure 3(a) shows a visualization of the data in a 2D-latent space using PCA projection, which preserves the geometric/periodic structure of the data. In this scenario, we expect GeoMatching’s advantage to be significant for those datapoints in the extremes of the manifold. Indeed, Figures 3(b) and 3(c) show that GeoMatching tends to match datapoints in the extremes to other datapoints inside the manifold, whereas Euclidean matching methods fail catastrophically, ignoring the manifold structure and matching several datapoints to the undesired outliers outside the manifold. All baselines are sensitive to the introduced outliers because they do not take into account data uncertainty; in contrast, GeoMatching is robust against the impact of outliers, which results in better ATE estimation, as shown in Figure 3(d). [^6] We note that Figure 3(d) looks less impressive than Figure 2(d) because ATE aggregates performance for all datapoints, and GeoMatching improves performance only for datapoints at the extremes of the manifold. These results align with our a priori expectations, as outliers tend to lie in low-density (high-volume) regions which are unlikely to be traversed by geodesics.
@@ 5.4 Real-world Causality Benchmark Datasets
We evaluate GeoMatching on two real-world datasets. We demonstrate that i) GeoMatching is competitive in more realistic scenarios, and ii) GeoMatching can benefit from semi-supervised settings, where we have access to several unlabelled covariate data, from which only a small subset has treatment and outcome information available. We consider the Infant Health Development Program Dataset and the Lalonde [^7] https://users.nber.org/~rdehejia/nswdata2.html dataset, see Appendix B for further details.
Table 1 summarizes the results on real-world datasets. For PCA, GeoMatching improves upon all baselines. We also include results for the Isomap projection. In this case, Euclidean matching in $\mathcal{Z}$ is the best approach, slightly above GeoMatching. We hypothesize that this is due to a coarse grid when optimizing the $\sigma$ parameter of the Riemannian metric (only 10 categorical values were explored in a grid search setup, see Appendix for details). In the case of the LaLonde dataset, using a Riemannian metric yield improvements in the Original covariate space, as well as in the PCA and NSM latent representations.